OpenStudy (anonymous):

Can someone explain how to get the derivative of 3x^2 + 15x + 5?

OpenStudy (anonymous):

there is a standard technique for taking the derivatives of a polynomial. first find the terms with similar degrees of x and combine them.

OpenStudy (anonymous):

(that's already done in this case, heh).

OpenStudy (anonymous):

Thanks guys I just started calculus

OpenStudy (anonymous):

then decrement the exponent by one, and multiply the former exponent with the coefficient.

OpenStudy (julie):

If you're deriving a single one of those terms, the move is this: \[ax^b\] The derivative of this is: \[a \cdot b \cdot x^{b - 1}\] The exception is for a constant (no power of x), which derives to 0.

OpenStudy (anonymous):

calculus is great! it opens up a whole new realm of math.

OpenStudy (julie):

In this case, you can do that for each one and you get: \[3 \cdot 2x + 15 \cdot 1 = 6x + 15\]

OpenStudy (anonymous):

looks like julie's got this one :)

OpenStudy (julie):

thanks for the medals :)